OpenStudy (anonymous):

Simplify. Please show all of the steps to arrive at your answer. -24 ÷ -2 ∙ -1/3

OpenStudy (anonymous):

-24 ÷ -2 = 24/2 = 12 so 12 . -1/3 = -12/3 = -4